OpenStudy (anonymous):

Using the formula A = lw, find the are of the rectangle \[9x ^{2}\] this is the length \[5x ^{2}\] this is the height

OpenStudy (anonymous):

\[area=length \times wide...or...area = length \times height\] you mean 5x^2 = wide ?

OpenStudy (anonymous):

\[5x^{2}\]

OpenStudy (anonymous):

is the height

OpenStudy (stamp):

Substitute in your length and width values. \[A=9x^2(5x^2)\]Evaluate. Properties of exponents may help you here:\[x^ax^b=x^{a+b}\]

OpenStudy (anonymous):

I think I got it \[45^{4}\]

OpenStudy (anonymous):

I multiplied the number and added the exponents

OpenStudy (stamp):

\[9x^2(5x^2)=9^15^1x^2x^2\]